10. Osmium is the densest metal with a density of 22.60 g/cc. (10 points) What is the volume, in cubic centimeters, of a block of osmium with dimensions 5.00 in X 4.00 in. X 0.25 ft?
Answer #1

10)

volume = 5.00 in * 4.00 in * 0.25 ft

volume = 5.00 in * 4.00 in * 0.25*12 in (because 1 ft = 12 in)

volume = 5.00 in * 4.00 in * 3.00 in

volume = 60.0 in^3

volume = 60.0 *(2.54 cm)^3 (because 1 inch = 2.5 cm)

volume = 983 cm^3

density = 22.60 g/cm^3

Answer: volume = 983 cm^3

------------------------------------------------------------------------------------------------------------------

If you need to calculate mass also, then only proceed further

so,

mass = density * volume

= 22.60 g/cm^3 * 983 cm^3

= 2.22*10^4 g